How much work does the electric field do in moving a proton from a point with a potential of +149 V to a point where it is -70 V? Express your answer both in joules and electron volts.
J
eV
Answer #1

|W| = 125 - (-55) = 180 eV

This converts to 2.88*10^-17 Joules
Answer #2
Work = (Charge)x(Voltage change)
The Charge on a proton is +e
W=(-70-149)=-219eV
1 electron volt = 1.60217646 × 10-19 joules
W=-219eV*1.60217646 × 10-19 joules
Answer #3
Charge on proton = 1.6 x 10 ^-19 C = 1e
Work in J = 1.6 x 10^-19 x(149-(-70)) =3.504 x10^-17 J
work in eV = 1 x (149+70) = 219 eV